Anti-differentiation Question 1. Find the function f(x) if 1. f(x) = 0.5e -0.2x, f(0) = 0 2. f(x) = x2 + Vx, f(1) = 3 3. f
Answer #1

1) -0.22 0.5 e flol D - flu) -0.2X dn Jo. 0.22 0.2 o floo D C 0:5 -0.5e + c 0.2 0.5 l - e-dian) 2.5 (1-e 0.2 -0.22 حمام - y f

flc - 3% => flu) lna te 7 Bio ५ dx = tc 4-1 4x-1 e Differentiating both sides + 4 4x7 e Чk 643-1 e Kat => Yk funt dut k in/4-

4 34+2 24 5k (3x+2) X 3 => 15k 1 k 2 15

1)

1. F(x) = 2.5(1-e-0.2x)

2. F(x) = x4/4+2/3x3/2 + 25/12

3. F(x) = ln(x) /7 +C

2)

1. K = - 1(There is negative sign in differentiation there will be - 4k=4)

2. K = - 1

3. K = 1/15

Hope you get your answer. Thank you and please like the answer.